OpenStudy (anonymous):

Which of the following is the correct solution of -2x - 3 ≥ 21? A. x ≥ -12 B. x ≤ 12 C. x ≤ -12 D. x ≥ 12

OpenStudy (tcp2315):

C, flip the sign when dividing or multiplying by a negative

OpenStudy (anonymous):

thanks guys :) i have 23 more anyone wanna help?? im so stressed and math is not what i want to be doing :(

OpenStudy (saifoo.khan):

\[-2x \ge 21 + 3\]\[-2x \ge 24\]\[0 \ge +2x +24\]FLIP\[2x + 24 \le 0\]

OpenStudy (saifoo.khan):

\[2x \le -24\]

OpenStudy (anonymous):

is a it C, OR D?

OpenStudy (tcp2315):

i havent done this in 4 years, im pretty sure its C, however i'll let him/her explain since they seem more familar with it, sorry if im wrong!

OpenStudy (saifoo.khan):

Sorry, i did a typo in the end.

OpenStudy (saifoo.khan):

TCP is correct.

OpenStudy (saifoo.khan):

\[x \le -12\] Rest of my working is correct.

OpenStudy (anonymous):

is a it C, OR D?

OpenStudy (anonymous):

c :)

OpenStudy (saifoo.khan):

C for Cat.

OpenStudy (tcp2315):

an honest mistake

OpenStudy (anonymous):

is a it C, OR D?

OpenStudy (anonymous):

no worries. thanks

OpenStudy (saifoo.khan):

hehe, right.

OpenStudy (anonymous):

is a it C, OR D?

OpenStudy (anonymous):

:D do you guys wanna help me with the assignment, lol i dont get math :(

OpenStudy (saifoo.khan):

It's C. C for Cat.

OpenStudy (saifoo.khan):

lol, which grade math is that.??

OpenStudy (tcp2315):

why dont you ask us question on how to do it, instead of just for the answers so when you do have a test you'll be able to do it yourself! i'll help

OpenStudy (anonymous):

is a it C, OR D?

OpenStudy (anonymous):

11th i think

OpenStudy (anonymous):

algebra II

OpenStudy (saifoo.khan):

it;s C. it;s C. it;s C. it;s C. it;s C. it;s C. it;s C. it;s C. it;s C.

OpenStudy (anonymous):

what is the solution of |x -10| =4 A. -6 or -14 B. -6 or 14 C. 6 or 14 D. 6 or -14

OpenStudy (anonymous):

??

OpenStudy (saifoo.khan):

C